Soln 1: 5 mL of unknown vit C diluted to 50 mL gives A = 0.358

Soln 2: 5 mL of unknown + 3 mL 12.3 mM vit C dil to 50 mL gives A = 0.871

Find the concentration of the unknown (hint: take ratio of absorbances)
Answer #1

5 ml of Unknown concetrtion has given absorbance at 0.358 Let keep this as X

5 ml of unknown concentration+3 ml of 12.3 mM std has given absorbance at 0.871

That x+y has given absorbance at 0.871

Hence Eq15X=0.358

Eq2 5X+3Y=0.871

Divide Eq1 by Eq2

5X/(5X+3Y)=0.358/0.871

do cross multiplication which give equation as follows

5X*(0.871)=5X+3Y*(0.358)

We know the concentration Y as 12.3 mM

5X*0.871=(5X+3*12.3)*0.358

4.355X=1.79X+13.2102

Rearrange equation as follows

4.355X-1.79X=13.2102

X(4.355-1.79)=13.2102

2.565X=13.2102

X=13.2102/2.565=5.150mM

Hence the conc of unknown is 5.150mM